OpenStudy (anonymous):

Find the component form of v given its magnitude and the angle it makes with the positive x-axis llvll=3√2 and θ= 150°

OpenStudy (anonymous):

\[ x= r\cos(\theta)\\ y= r\sin(\theta)\\ \]

OpenStudy (anonymous):

\[ \cos (150 {}^{\circ})=-\frac{\sqrt{3}}{ 2}\\ \cos (150 {}^{\circ})=\frac{1}{ 2} \]

OpenStudy (anonymous):

\[ x=-\frac{3}{2} \sqrt{2} \sqrt{3}=-3 \sqrt{\frac{3}{2}} \]

OpenStudy (anonymous):

\[ y=\frac{3 \sqrt{2}}{2}=\frac{3}{\sqrt{2 }} \]